Shift Sum: DAY Nov 5, 1999

Douglas A. Jensen (
Fri, 5 Nov 1999 15:47:33 -0600 (CST)

Shift Summary for DAY shift, Nov 5, 1999

On Shift: Jensen, Wang

Hours of beam delivered at all intensities: 6
Hours of beam delivered at nominal intensity: 1
Hours of 99 data (special runs not included): 2.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
Repairs in the hall, interlock reset, beam off for SWIC check,
during which time check RC8


Runs ended this shift:
Run 14906

Tapes brought to FCC this shift: none


Overview of shift:
Ran for debugging, up and down during the am. Then at noon,
requested full intensity, continuing the debugging run. Up
to full intensity around 1 pm.

Description of Shift:

run spill time comment
14905 10:55 End. Debugging run
14910 11:05 Checking, intensity increase...

13:15 Get TRD alarm. Paged experts.
13:30 L2BEAM/L!BEAM error message. ????
13:45 beam will go off for a while - Rick Ford et al
need to check a SWIC. Sasha returned page,
reported TRD problem. He will check. Elizabeth
will make an access and work on RC8.
15:00 Rick Ford calls. The found a SWIC cable
unplugged. Also found a hot spot upstream, so
beam was hitting something. Thinking....

14911 15:10 Run started. One CSI constant changed.
15:15 KTeV done with access. MCR had called to check
a bit earlier.
15:30 MCR says that they are turning on. But no beam
yet. Shift ends....
Follow up Notes on Previous Shift
HOT/funny HCC bit fixed. Acc. scaler now working, L2/L1 rates
much more reasonable. Expect RC8 fixed. SWIC fixed.

Items to Follow up on Next Shift
Take data !